Express each number as a product of its prime factors:
3825

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

3825
By Taking the LCM of 3825, we will get the product of its prime factor.
Hence,3825 = 3 × 3 × 5 × 5 × 17 × 1 = 3^2×5^2×17

​Thanks
Ram Kushwah
3 years ago
3825/3=1275
1275/3=425
425/5=85
85/5=17
17 is  prime number
so the number 3825 can be written as :
3825=3x3x5x5x17
=3² x 5² x 17